The weight of potato chips in a medium‐size bag is stated to be 10 ounces. The amount that the packaging machine puts in these bags is believed to have a Normal distribution with a mean of 10.2 ounces and a standard deviation of 0.14 ounces.
1.If the production engineer wants to change the value of the mean but maintain the standard deviation at 0.14 ounces so that at most 1.5% of all bags are underweight, how large does the engineer need to make that mean?
1
Expert's answer
2019-11-13T11:58:38-0500

Let "X=" the weight of potato chips in ounces: "X\sim N(\mu, \sigma^2)"

Then


"Z={X-\mu\over \sigma}\sim N(0,1)"

Given that "\mu_0=10.2 \ ounces, \sigma=0.14\ ounces"

If "P(X<10)=0.015," find "\mu_{new}"


"P(X<10)=P(Z<{10-\mu_{new}\over 0.14})=0.015"

Then


"{10-\mu_{new}\over 0.14}\approx-2.170090"

"\mu_{new}=10+0.14(2.170090)\approx10.3"

"\mu_{new}-\mu_0=10.3-10.2=0.1"

The  engineer needs to increase the mean by "0.1" ounces. New mean will be "10.3" ounces.